# AudioMoG: Guiding Audio Generation with Mixture-of-Guidance

### v1

@@ F.3 Model configurations
Our diffusion model is built upon the Diffusion Transformer (DiT) (Peebles & Xie, 2023) architecture, following a latent diffusion modeling (LDM) (Rombach et al., 2022) paradigm that offers strong generative capabilities and effective context modeling. The backbone of the diffusion network adopts a DiT structure with 24 layers and 24 attention heads, each with an embedding dimension of 1536. The model supports both cross-attention and global conditioning: cross-attention is applied to all types of conditional inputs, while global conditioning specifically handles duration-related control signals. The internal token dimension of the diffusion model is set to 64, with a conditional token dimension of 768 and a global condition embedding dimension of 1536. The generated latent representation has the same dimensionality as io_channels, which is 64.
F.4 Compression networks
@@
To train the audio autoencoder, we adopt a variational autoencoder (VAE) (Kingma et al., 2013) architecture based on the Oobleck framework (Evans et al., 2025) at a sampling rate of 16kHz. The model is trained from scratch on large-scale publicly available text-audio paired datasets. The encoder and decoder are symmetric, each using a base channel size of 128, with channel multipliers $1,2,4,8,16$ and strides $2,2,4,4,10$. The encoder maps the input waveform into a 128-dimensional latent representation, while the decoder reconstructs the waveform from a 64-dimensional latent code. Snake activation is applied throughout the network, and no final tanh activation is used in the decoder. The overall downsampling ratio is 640, and both input and output are mono-channel waveforms. The bottleneck is implemented as a variational layer.
F.5 Objective metrics
@@
We introduce the objective metrics employed in our evaluation, including Fréchet Audio Distance (FAD) (Kilgour et al., 2019), Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951), Inception Score (IS) (Salimans et al., 2016), Fréchet Distance (FD) (Heusel et al., 2017), and LAION-CLAP score (Wu et al., 2023). FAD, adapted from FID (Heusel et al., 2017), measures the distributional gap between generated and reference audio using VGGish embeddings (Hershey et al., 2017), and serves as our primary indicator of audio fidelity. KL divergence evaluates the difference in acoustic event posteriors between ground truth and generated audio, computed using the PANN tagging model (Kong et al., 2020). IS reflects both diversity and specificity of the generated samples, based on entropy over class predictions. FD, while similar in formulation to FAD, is computed in more general embedding spaces and tends to be less stable in audio tasks. We include it for completeness but primarily rely on FAD for fidelity assessment. The CLAP score is calculated as the cosine similarity between CLAP embeddings of the generated audio and the corresponding text. We use the AudioLDM evaluation toolkit to compute all objective metrics.
F.6 Subjective evaluation
@@
We randomly selected 20 samples from the AudioCaps test set for the subjective evaluation. Each group includes the results from AudioLDM, AudioLDM2, CFG-only, HG, and the ground truth (GT), with the order of samples within each group randomly shuffled. Each group was rated by 20 human raters. In our evaluation, both overall quality (OVL) and text relevance (REL) are rated on a scale from 1 to 5. For OVL, raters assess the perceptual quality of the audio, while for REL, they rate the relevance of the audio to the given text condition. The minimum rating increment for all scores is 1 point. A screenshot of our evaluation interface is shown in Figure 7.
@@ Appendix G Video-to-audio experiment details
G.1 Datasets
Apart from T2A and T2M generation, we conduct experiments for video-to-audio(V2A) generation. We utilize the benchmark datasets AudioSet (Gemmeke et al., 2017) and VGGSound (Chen et al., 2020) for model training. To compare with previous work, we evaluated our models on the VGGSound test set, which consists of about 15K 10-second audio clips.
G.2 Model configurations
We use CLIP (Radford et al., 2021) embeddings to extract the visual features, and we leverage the same VAE as in T2A. The diffusion backbone shares the same core architecture as the T2A counterpart, being built upon the DiT within the LDM paradigm. Most hyperparameters remain the same with T2A, but we increased the conditional token dimension to 1024 to better accommodate high-dimensional visual embeddings, and the global condition embedding dimension is set to 2048.
We fine-tuned from a T2A model that was trained for 2M iterations with a batch size of 8 per GPU. For the good model, we conducted 1.3M finetuning iterations, while for the bad model, we fine-tuned for 0.3M iterations, both using a batch size of 8 per GPU. The optimizer and sampler settings are the same as those used in the T2A model.
G.3 Baselines methods
@@
IM2WAV is an image-guided open-domain audio generation system introduced by (Sheffer & Adi, 2023). It employs two Transformer language models operating over a discrete audio representation derived from a VQ-VAE (Van Den Oord et al., 2017) model. The system first generates a low-level audio representation using a language model, then upsamples the audio tokens with an additional language model to produce high-fidelity audio. Visual conditioning is achieved through CLIP embeddings, and CFG is applied to steer the generation process.
@@
Diff-Foley is a synchronized V2A synthesis method utilizing a latent diffusion model (LDM), presented by (Luo et al., 2023). It incorporates contrastive audio-visual pretraining (CAVP) to learn temporally and semantically aligned features, which are then used to train the LDM on spectrogram latent space. The model employs cross-attention modules and "double guidance" to enhance sample quality, achieving state-of-the-art performance in V2A tasks.
@@
FoleyGen is an open-domain V2A generation system based on a language modeling paradigm, introduced by (Mei et al., 2024). It leverages a neural audio codec for bidirectional conversion between waveforms and discrete tokens. A single Transformer model, conditioned on visual features extracted from a visual encoder, facilitates the generation of audio tokens. The model addresses temporal synchronization challenges by exploring novel visual attention mechanisms.
@@
VTA-LDM is a V2A generation framework developed by (Xu et al., 2024), building upon the LDM framework. It employs a CLIP-based vision encoder to extract frame-level video features, which are temporally concatenated and mapped using a projector as the generation condition. The model focuses on generating semantically and temporally aligned audio content corresponding to video inputs.
@@
FoleyCrafter (Zhang et al., 2024) is a text-based V2A generation framework designed to produce high-quality, semantically relevant, and temporally synchronized audio for videos. It extends state-of-the-art T2A generators by incorporating a semantic adapter for semantic alignment and a temporal adapter for precise audio-video synchronization, ensuring realistic sound effects that align with visual content.
@@
V2A-Mapper is a lightweight solution for V2A generation proposed by (Wang et al., 2024a). It connects foundation models by translating visual CLIP embeddings into auditory CLAP embeddings, bridging the domain gap between visual and audio modalities. Conditioned on the translated CLAP embedding, a pretrained audio generative model (AudioLDM) is used to produce high-fidelity and visually-aligned sound, requiring minimal training parameters.
@@
VAB-Encodec (Su et al., 2024) is a unified audio-visual framework that learns latent representations and enables vision-to-audio generation within the same model. It tokenizes 48 kHz audio with a pretrained Encodec tokenizer and encodes video frames with an image encoder. During pre-training the model performs visual-conditioned masked-audio-token prediction; at inference it iteratively decodes audio tokens conditioned on visual features, yielding fast and semantically aligned sound.
@@
VATT is a multi-modal generative framework for V2A generation through text, presented by (Liu et al., 2024b). It comprises two modules: VATT Converter, a large language model fine-tuned for instructions that maps video features to the LLM vector space; and VATT Audio, a transformer that generates audio tokens from visual frames and optional text prompts using iterative parallel decoding. The framework allows for controllable audio generation and audio captioning based on video inputs.
@@
For Diff-foley, VTA-LDM and FoleyCrafter, we generate 10-second audio samples using their official implementations. For V2A Mapper, it supplies pre-generated audio samples for evaluation. As the official implementations of FoleyGen, VAB-Encodec and VATT are unavailable, we compare our results with their official reported results. The IM2WAV results are adopted from VATT.
G.4 Metrics
@@
We introduce additional metrics utilized in our V2A evaluation apart from FAD, KL, IS and FD, including Imagebind Score (IBS) and temporal alignment accuracy (Align acc), which primarily measures audio-visual alignment.
@@
ImageBind Score (IBS) assesses the semantic alignment between generated audio and the corresponding video by computing the cosine similarity between their embeddings in a shared multimodal space. This metric leverages the ImageBind model (Girdhar et al., 2023), which aligns multiple modalities—including images, audio, and text—into a unified embedding space, facilitating cross-modal retrieval and evaluation . A higher IBS indicates a stronger semantic correlation between the audio and video content.
@@
Temporal Alignment Accuracy (Align Acc) measures the synchronization between generated audio and video by evaluating the model’s ability to produce audio events that are temporally aligned with visual events. Introduced in Diff-Foley (Luo et al., 2023), this metric involves training a classifier to distinguish between correctly aligned audio-video pairs and misaligned ones. The classifier is trained on three types of pairs: true pairs (correctly aligned), temporally shifted pairs, and mismatched pairs from different videos. Align Acc is computed as the percentage of correctly identified true pairs, providing a quantitative measure of temporal synchronization.
@@
Appendix H Detailed related work
H.1 Text-to-audio (T2A) generation
T2A systems generate audio samples conditioned on natural language prompts. At the beginning, DiffSound (Yang et al., 2023), AudioGen (Kreuk et al., 2022) explore autoregressive-based generation methods in the compressed space of mel-spectrogram and waveform respectively. Then, AudioLDM (Liu et al., 2023) and Make-An-Audio (Huang et al., 2023b) develop latent diffusion models in the compressed space of mel-spectrogram, improving overall T2A generation quality. Tango (Deepanway et al., 2023) improves the text encoder of diffusion-based T2A systems with a language model. AudioLDM 2 (Liu et al., 2024a) employs an autoregressive-based method to predict the AudioMAE (Huang et al., 2022) features from various input modalities, and then uses a latent diffusion model to generate audio from AudioMAE features. Tango2 (Majumder et al., 2024) and Tangoflux (Hung et al., 2024) utilize reinforcement strategies including direct preference optimization (Rafailov et al., 2023) and CLAP-guided reward shaping (Wu et al., 2023) to improve the human preference and semantic-textual alignment. Recently, Stable Audio (Evans et al., 2024) designs transformer-based scalable latent diffusion models in the space directly compressed from the audio waveform. ETTA (Lee et al., 2025) elucidates the design space of diffusion-based T2A systems.
@@ H.1 Text-to-audio (T2A) generation
H.2 Video-to-audio (V2A) generation
@@
Recent advances in video-to-audio (V2A) generation can be broadly divided into two categories: (1) enhancing V2A via pre-trained text-to-audio (T2A) models, and (2) introducing auxiliary temporal representations to improve temporal alignment. In the first category, methods such as V2A-Mapper (Wang et al., 2024a) and FoleyCrafter (Mei et al., 2024) build upon established T2A systems like AudioLDM (Liu et al., 2023). These approaches either align video features with the original conditioning space of T2A models or introduce additional adapters to inject visual information as supplementary conditions. For example, V2A-Mapper proposes a mapping strategy that translates video features into audio CLAP embeddings, enabling AudioLDM to perform V2A synthesis. Similarly, FoleyCrafter integrates dedicated adapters to incorporate visual cues into the conditioning process of T2A models.
@@
The second category includes methods such as TiVA (Wang et al., 2024b), ReWaS (Jeong et al., 2024), SyncFusion (Comunità et al., 2024), and SonicVisionLM (Xie et al., 2024), which incorporate explicitly designed temporal features to enhance synchronization between video and audio. TiVA employs downsampled Mel spectrograms as auxiliary representations that carry temporal structure, and utilizes a transformer-based predictor to estimate these features for guiding V2A generation. SyncFusion and SonicVisionLM leverage onset positions and audio timestamps, respectively, as temporal control signals during synthesis. ReWaS introduces energy as a continuous temporal representation, providing a more fine-grained condition along the time axis to better regulate V2A output. At the sampling stage, CFG is popularly employed in these methods. In our paper, we explore a novel sampling algorithm to increase generation quality in a training-free and computationally lightweight manner, which is orthogonal to previous innovations.
H.3 Guidance methods
@@
CFG. In previous diffusion-based T2A generation  (Liu et al., 2023; 2024a; Huang et al., 2023b; a; Evans et al., 2024; 2025; Lee et al., 2025) and V2A generation works (Sheffer & Adi, 2023; Luo et al., 2023; Mei et al., 2024; Xu et al., 2024; Wang et al., 2024a; Su et al., 2024; Liu et al., 2024b) , CFG (Ho & Salimans, 2022) is commonly adopted to improve the audio generation quality at the inference stage. To achieve optimal results, its guidance scale is investigated in different methods (Hung et al., 2024). However, as demonstrated in recent work (Lee et al., 2025), CFG sacrifices the diversity of generation results and may suffer from suboptimal synthesis quality. Previous theoretical analysis (Chidambaram et al., 2024) demonstrates that for any non-zero level of score estimation error, a large CFG strength causes the sampler to diverge from the data distribution’s support, formally explaining the empirical phenomenon of distortion at high guidance scales. Our analysis reveals that when CFG uses a bad unconditional model, it inherently introduces a score correction term. This may explain the empirical finding that a small unconditional model can effectively guide a large conditional model (Karras et al., 2024b). However, in standard CFG, this beneficial correction term is entangled with the conditional alignment term, making it difficult to find a guidance strength that simultaneously ensures outlier removal and quality enhancement.
@@
Guidance with weak models. Recently, AG (Karras et al., 2024a) proposes a method to guide a diffusion model with the bad version of itself, demonstrating stronger synthesis quality than CFG in image domain. Several works (Phunyaphibarn et al., 2025; Jeon, 2025; Hyung et al., 2025) has extended this idea to the scenarios of fine-tuning diffusion models to a specific task (Phunyaphibarn et al., 2025), motion synthesis (Jeon, 2025), and video generation (Hyung et al., 2025). Other works adopt similar ideas to construct a weak model. For instance, SAG (Hong et al., 2023) applies Gaussian blurring to the model input, SG (Li et al., 2025) alters the denoising timestep to obtain an output with higher noise levels, while ICG (Sadat et al., 2024) randomly samples a condition to replace the null embedding in CFG. Among these guidance methods, AG is the most representative. It provides a theoretical justification for its efficacy, positing that it reduces the score estimation error induced by the "mass-covering" or "mean-seeking" behavior of the score matching training objective. However, the advantages of AG have not been observed for audio generation. As recently mentioned in ETTA (Lee et al., 2025), AG is sensitive to the choice of the bad model, prohibiting their application on audio synthesis. In this work, we explore the advantages of AG for audio generation, and propose a novel sampling algorithm, MoG, yielding cumulative advantages of AG and CFG to outperform either of them on audio synthesis.

### v2

@@ F-C Model configurations
Our diffusion model is built upon the Diffusion Transformer (DiT) architecture, following a latent diffusion modeling (LDM) paradigm that offers strong generative capabilities and effective context modeling. We use FLAN-T5 as the text encoder for our model, and we train a Variational Autoencoder (VAE) that compresses the original waveform into the latent representation. The backbone of the diffusion network adopts a DiT structure with 24 layers and 24 attention heads, each with an embedding dimension of 1536. The model supports both cross-attention and global conditioning: cross-attention is applied to all types of conditional inputs, while global conditioning specifically handles duration-related control signals. The internal token dimension of the diffusion model is set to 64, with a conditional token dimension of 768 and a global condition embedding dimension of 1536. The generated latent representation has the same dimensionality as io_channels, which is 64.
F-D Compression networks
@@
To train the audio autoencoder, we adopt a variational autoencoder (VAE) architecture based on the Oobleck framework [14] at a sampling rate of 16kHz. The model is trained from scratch on large-scale publicly available text-audio paired datasets. The encoder and decoder are symmetric, each using a base channel size of 128, with channel multipliers $1,2,4,8,16$ and strides $2,2,4,4,10$. The encoder maps the input waveform into a 128-dimensional latent representation, while the decoder reconstructs the waveform from a 64-dimensional latent code. Snake activation is applied throughout the network, and no final tanh activation is used in the decoder. The overall downsampling ratio is 640, and both input and output are mono-channel waveforms. The bottleneck is implemented as a variational layer.
F-E Objective metrics
@@
We introduce the objective metrics employed in our evaluation, including Fréchet Audio Distance (FAD), Kullback-Leibler (KL) divergence, Inception Score (IS), Fréchet Distance (FD), and LAION-CLAP score [57]. FAD, adapted from FID (Frechet Inception Distance), measures the distributional gap between generated and reference audio using VGGish embeddings, and serves as our primary indicator of audio fidelity. KL divergence evaluates the difference in acoustic event posteriors between ground truth and generated audio. IS reflects both diversity and specificity of the generated samples, based on entropy over class predictions. FD, while similar in formulation to FAD, is computed in more general embedding spaces and tends to be less stable in audio tasks. We include it for completeness but primarily rely on FAD for fidelity assessment. The CLAP score is calculated as the cosine similarity between CLAP embeddings of the generated audio and the corresponding text. We use the AudioLDM evaluation toolkit to compute all objective metrics.
F-F Subjective evaluation
@@
We randomly selected 20 samples from the AudioCaps test set for the subjective evaluation. Each group includes the results from AudioLDM, AudioLDM2, CFG-only, HG, and the ground truth (GT), with the order of samples within each group randomly shuffled. Each group was rated by 20 human raters. In our evaluation, both overall quality (OVL) and text relevance (REL) are rated on a scale from 1 to 5. For OVL, raters assess the perceptual quality of the audio, while for REL, they rate the relevance of the audio to the given text condition. The minimum rating increment for all scores is 1 point. A screenshot of our evaluation interface is shown in Fig. 6.
@@ Appendix G Video-to-audio experiment details
G-A Datasets
Apart from T2A and T2M generation, we conduct experiments for video-to-audio (V2A) generation. We utilize the benchmark datasets AudioSet [17] and VGGSound [4] for model training. To compare with previous work, we evaluated our models on the VGGSound test set, which consists of about 15K 10-second audio clips.
G-B Model configurations
We use CLIP embeddings to extract the visual features, and we leverage the same VAE as in T2A. The diffusion backbone shares the same core architecture as the T2A counterpart, being built upon the DiT within the LDM paradigm. Most hyperparameters remain the same with T2A, but we increased the conditional token dimension to 1024 to better accommodate high-dimensional visual embeddings, and the global condition embedding dimension is set to 2048.
We fine-tuned from a T2A model that was trained for 2M iterations with a batch size of 8 per GPU. For the good model, we conducted 1.3M finetuning iterations, while for the bad model, we fine-tuned for 0.3M iterations, both using a batch size of 8 per GPU. The optimizer and sampler settings are the same as those used in the T2A model. In the main results Table III, we set $w_{1}=2.7,w_{2}=0.15$ in MoG-PG and $w_{1}=2,7,w_{2}=2.5,w_{3}=1.2$ in MoG-HG.
G-C Baselines methods
@@
IM2WAV is an image-guided open-domain audio generation system introduced by [53]. It employs two Transformer language models operating over a discrete audio representation derived from a VQ-VAE model. The system first generates a low-level audio representation using a language model, then upsamples the audio tokens with an additional language model to produce high-fidelity audio. Visual conditioning is achieved through CLIP embeddings, and CFG is applied to steer the generation process.
@@
Diff-Foley is a synchronized V2A synthesis method utilizing a latent diffusion model (LDM), presented by [46]. It incorporates contrastive audio-visual pretraining (CAVP) to learn temporally and semantically aligned features, which are then used to train the LDM on spectrogram latent space. The model employs cross-attention modules and "double guidance" to enhance sample quality, achieving state-of-the-art performance in V2A tasks.
@@
FoleyGen is an open-domain V2A generation system based on a language modeling paradigm, introduced by [48]. It leverages a neural audio codec for bidirectional conversion between waveforms and discrete tokens. A single Transformer model, conditioned on visual features extracted from a visual encoder, facilitates the generation of audio tokens. The model addresses temporal synchronization challenges by exploring novel visual attention mechanisms.
@@
VTA-LDM is a V2A generation framework developed by [59], building upon the LDM framework. It employs a CLIP-based vision encoder to extract frame-level video features, which are temporally concatenated and mapped using a projector as the generation condition. The model focuses on generating semantically and temporally aligned audio content corresponding to video inputs.
@@
FoleyCrafter [61] is a text-based V2A generation framework designed to produce high-quality, semantically relevant, and temporally synchronized audio for videos. It extends state-of-the-art T2A generators by incorporating a semantic adapter for semantic alignment and a temporal adapter for precise audio-video synchronization, ensuring realistic sound effects that align with visual content.
@@
V2A-Mapper is a lightweight solution for V2A generation proposed by [55]. It connects foundation models by translating visual CLIP embeddings into auditory CLAP embeddings, bridging the domain gap between visual and audio modalities. Conditioned on the translated CLAP embedding, a pretrained audio generative model (AudioLDM) is used to produce high-fidelity and visually-aligned sound, requiring minimal training parameters.
@@
VAB-Encodec [54] is a unified audio-visual framework that learns latent representations and enables vision-to-audio generation within the same model. It tokenizes 48 kHz audio with a pretrained Encodec tokenizer and encodes video frames with an image encoder. During pre-training the model performs visual-conditioned masked-audio-token prediction; at inference it iteratively decodes audio tokens conditioned on visual features, yielding fast and semantically aligned sound.
@@
VATT is a multi-modal generative framework for V2A generation through text, presented by [44]. It comprises two modules: VATT Converter, a large language model fine-tuned for instructions that maps video features to the LLM vector space; and VATT Audio, a transformer that generates audio tokens from visual frames and optional text prompts using iterative parallel decoding. The framework allows for controllable audio generation and audio captioning based on video inputs.
@@
For Diff-foley, VTA-LDM and FoleyCrafter, we generate 10-second audio samples using their official implementations. For V2A-Mapper, it supplies pre-generated audio samples for evaluation. As the official implementations of FoleyGen, VAB-Encodec and VATT are unavailable, we compare our results with their official reported results. The IM2WAV results are adopted from VATT.
G-D Metrics
@@
We introduce additional metrics utilized in our V2A evaluation apart from FAD, KL, IS and FD, including Imagebind Score (IBS) and temporal alignment accuracy (AA), which primarily measures audio-visual alignment.
@@
ImageBind Score (IBS) assesses the semantic alignment between generated audio and the corresponding video by computing the cosine similarity between their embeddings in a shared multimodal space. This metric leverages the ImageBind model [19], which aligns multiple modalities—including images, audio, and text—into a unified embedding space, facilitating cross-modal retrieval and evaluation . A higher IBS indicates a stronger semantic correlation between the audio and video content.
@@
Temporal Alignment Accuracy (AA) measures the synchronization between generated audio and video by evaluating the model’s ability to produce audio events that are temporally aligned with visual events. Introduced in Diff-Foley [46], this metric involves training a classifier to distinguish between correctly aligned audio-video pairs and misaligned ones. The classifier is trained on three types of pairs: true pairs (correctly aligned), temporally shifted pairs, and mismatched pairs from different videos. Align Acc is computed as the percentage of correctly identified true pairs, providing a quantitative measure of temporal synchronization.
@@
Appendix H Detailed related works
H-A Text-to-audio (T2A) generation
T2A systems generate audio samples conditioned on natural language prompts. At the beginning, AudioGen [35] explore autoregressive-based generation methods in the compressed space of mwaveform respectively. Then, DiffSound [60], AudioLDM [42] and Make-An-Audio [23] develop latent diffusion models in the compressed space of mel-spectrogram, improving overall T2A generation quality. Tango [18] improves the text encoder of diffusion-based T2A systems with a language model. AudioLDM 2 [43] employs an autoregressive-based method to predict the AudioMAE features from various input modalities, and then uses a latent diffusion model to generate audio from AudioMAE features. Tango2 [47] and Tangoflux [24] utilize reinforcement strategies including CLAP-guided reward shaping [57] to improve the human preference and semantic-textual alignment. Recently, Stable Audio [13] designs transformer-based scalable latent diffusion models in the space directly compressed from the audio waveform. ETTA [38] elucidates the design space of diffusion-based T2A systems.
@@ H-A Text-to-audio (T2A) generation
H-B Video-to-audio (V2A) generation
@@
Recent advances in video-to-audio (V2A) generation can be broadly divided into two categories: (1) enhancing V2A via pre-trained text-to-audio (T2A) models, and (2) introducing auxiliary temporal representations to improve temporal alignment. In the first category, methods such as V2A-Mapper [55] and FoleyCrafter [48] build upon established T2A systems like AudioLDM [42]. These approaches either align video features with the original conditioning space of T2A models or introduce additional adapters to inject visual information as supplementary conditions. For example, V2A-Mapper proposes a mapping strategy that translates video features into audio CLAP embeddings, enabling AudioLDM to perform V2A synthesis. Similarly, FoleyCrafter integrates dedicated adapters to incorporate visual cues into the conditioning process of T2A models.
@@
The second category includes methods such as TiVA [56], ReWaS [27], SyncFusion [6], and SonicVisionLM [58], which incorporate explicitly designed temporal features to enhance synchronization between video and audio. TiVA employs downsampled Mel spectrograms as auxiliary representations that carry temporal structure, and utilizes a transformer-based predictor to estimate these features for guiding V2A generation. SyncFusion and SonicVisionLM leverage onset positions and audio timestamps, respectively, as temporal control signals during synthesis. ReWaS introduces energy as a continuous temporal representation, providing a more fine-grained condition along the time axis to better regulate V2A output. At the sampling stage, CFG is popularly employed in these methods. In our paper, we explore a novel sampling algorithm to increase generation quality in a training-free and computationally lightweight manner, which is orthogonal to previous innovations.
H-C Guidance methods
@@
CFG. In previous diffusion-based T2A generation [42, 43, 23, 22, 13, 14, 38, 29, 28, 8] and V2A generation works [53, 46, 48, 59, 55, 54, 44, 9] , CFG [20] is commonly adopted to improve the audio generation quality at the inference stage. To achieve optimal results, its guidance scale is investigated in different methods [24]. However, as demonstrated in recent work [38], CFG sacrifices the diversity of generation results and may suffer from suboptimal synthesis quality. Previous theoretical analysis [5] demonstrates that for any non-zero level of score estimation error, a large CFG strength causes the sampler to diverge from the data distribution’s support, formally explaining the empirical phenomenon of distortion at high guidance scales. Our analysis reveals that when CFG uses a bad unconditional model, it inherently introduces a score correction term. This may explain the empirical finding that a small unconditional model can effectively guide a large conditional model [30]. However, in standard CFG, this beneficial correction term is entangled with the conditional alignment term, making it difficult to find a guidance strength that simultaneously ensures outlier removal and quality enhancement.
@@
Guidance with weak models. Recently, AG [31] proposes a method to guide a diffusion model with the bad version of itself, demonstrating stronger synthesis quality than CFG in image domain. Several works [50, 26, 25] has extended this idea to the scenarios of fine-tuning diffusion models to a specific task [50], motion synthesis [26], and video generation [25]. Other works adopt similar ideas to construct a weak model. For instance, SAG [21] applies Gaussian blurring to the model input, SG [41] alters the denoising timestep to obtain an output with higher noise levels, while ICG [51] randomly samples a condition to replace the null embedding in CFG. Among these guidance methods, AG is the most representative. It provides a theoretical justification for its efficacy, positing that it reduces the score estimation error induced by the "mass-covering" or "mean-seeking" behavior of the score matching training objective. However, the advantages of AG have not been observed for audio generation. As recently mentioned in ETTA [38], AG is sensitive to the choice of the bad model, prohibiting their application on audio synthesis. In this work, we explore the advantages of AG for audio generation, and propose a novel sampling algorithm, MoG, yielding cumulative advantages of AG and CFG to outperform either of them on audio synthesis.